# The Relative Entropy of Expectation and Price

### v1

@@ III Entropic pricing and hedging
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The case studies in Figures 6 and 7 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer spread around the mid model that prices the residual risk after hedging. To further develop intuition on the meaning of the hedge and price conditions and their solutions, it is useful to consider boundary cases where they simplify.
@@ IV Market and model risk
Continuing from the previous section, with the expectation measure $\mathbb{E}$ the calibration conditions:
@@ V Entropic XVA
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Returning to the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a transaction with initial price $p$, where $-c_{2}\leq p\leq c_{1}$. Excess capital is deposited in the margin account with price $b_{1}>0$ and $b_{2}>0$ respectively. The transaction is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.

### v2

@@ III Entropic pricing and hedging
$\rho=-0.5$ $\rho=-0.5$
Refer to caption Refer to caption
$\rho=0$ $\rho=0$
Refer to caption Refer to caption
$\rho=0.5$ $\rho=0.5$
Figure 6: Discounting is an emergent feature in entropic pricing, where the funding rate is implied as the Lagrange multiplier that implements the self-funding condition. In this example, two underlying securities are modelled with lognormal final prices $q+dq=q(1+\mu\,dt)\exp[\sigma\sqrt{dt}\,z-\sigma^{2}dt/2]$ in the expectation measure, using standard normal variables $z$ that have correlation $\rho$. The model for the second underlying security is fixed with $q_{2}=1$, $\mu_{2}=5\%$ and $\sigma_{2}=20\%$. The model for the first underlying security uses $q_{1}=1$, and the graphs in each row plot the implied funding rate $r$ and unit optimal portfolio $\phi_{1}$ as $\mu_{1}$ and $\sigma_{1}$ are varied. The correlation is $\rho=-0.5$ in the first row, $\rho=0$ in the second row, and $\rho=0.5$ in the third row, and the investment horizon is set to $dt=1$.
      When the volatility $\sigma_{1}$ of the first underlying security is small, the implied funding rate $r$ aligns with the mean rate $\mu_{1}$, as in this case the underlying security approximates a funding security with predictable return. When the volatilities $\sigma_{1}$ and $\sigma_{2}$ of the underlying securities are close to each other, the implied funding rate $r$ is approximately the average of the two mean rates $\mu_{1}$ and $\mu_{2}$. The mean rates of the underlying securities thus contribute to the implied funding rate, which is essentially the predictability-weighted average of the mean rates.
Refer to caption Refer to caption
@@
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The case studies in Figures 7 and 8 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer spread around the mid model that prices the residual risk after hedging. To further develop intuition on the meaning of the hedge and price conditions and their solutions, it is useful to consider boundary cases where they simplify.
@@ III.5 Heston model
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimal entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk adjustment is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimal entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk adjustment is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@ IV Market and model risk
Returning to the general pricing framework, with the expectation measure $\mathbb{E}$ the calibration conditions:
@@ V Entropic XVA
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Revisiting the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a transaction with initial price $p$, where $-c_{2}\leq p\leq c_{1}$. Excess capital is deposited in the margin account with price $b_{1}>0$ and $b_{2}>0$ respectively. The transaction is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.